# Minimax surfaces and the holographic entropy cone

### v1

@@ 1 Introduction
For multiple regions, we define the notion of a cooperating time-sheet configuration, as follows: Intersecting time-sheets cut each other up into partial time-sheets, and cut the surfaces they host into partial surfaces; a set of minimax time-sheets cooperates if every partial HRT surface is maximal on its partial time-sheet. The value of cooperating time-sheets is that, as we show, they define a graph model. We show that, for a pair of crossing regions with connected HRT surfaces, there does exist a cooperating time-sheet pair. However, we have not been able to prove that a cooperating configuration exists for a general set of boundary regions, and our investigations of specific examples are inconclusive; while in some cases we can construct cooperating time-sheets, in others we can actually argue against their existence. This leaves several logically possible outcomes for this story:
@@ 1.1 User-friendly summary
In this section we provide an overview of the main results; proofs and details are given in the main text. In what follows, the bulk spacetime will be denoted by $\hat{\mathcal{M}}$, the conformal boundary by $\hat{\mathcal{N}}$, and the future and past boundaries by $\hat{\mathcal{I}}^{+}$ and $\hat{\mathcal{I}}^{-}$ respectively (we use the hatted notation to avoid confusion with the regulated spacetime that we will use in the bulk of the paper, which will be introduced in section 2.1). For a fixed boundary region $A$ on a boundary Cauchy slice $\Sigma\subset\hat{\mathcal{N}}$, the conformal boundary can then be decomposed as (cf. figure 1)
@@
Except for the case of two intersecting time-sheets, we do not have a general proof of the existence of cooperating time-sheets, which we leave as a conjecture. However, we show supporting evidence of its validity in subsection 4.2 by doing direct analytic and numerical computations for a specific set of configurations of time-sheets in pure AdS3 and in asymptotically AdS3 with spherically symmetric NEC-obeying matter. In subsection 4.3, we also present a scenario in pure AdS3 where the existence of a cooperating configuration for three time-sheets is challenged. The setup entails an apparently frustrated configuration of nearly-intersecting but relatively-boosted HRT surfaces. In this extreme scenario cooperation appears difficult to achieve, suggesting that either a more clever construction is needed or the cooperating conjecture, as currently stated, may be false. We leave further investigations to future studies.
@@ 2.3 Stability and extremality of minimax surfaces
Of course, these theorems do not preclude a region $A$ having multiple, degenerate minimax/HRT surfaces. However, these surfaces must be maximal on different time-sheets. In subsec. 3.3 below, we will prove a stronger result: degenerate minimax surfaces must be fully spacelike-separated. The same result is also obtained from maximin, where it is shown that degenerate HRT surfaces share a common Cauchy slice Wall_2014 ; Grado-White:2024gtx .
@@ 2.4 Relaxed minimax
As noted in subsec. 2.2, the definition of time-sheet implies that $\tau$ intersects $\gamma^{0}$ (if the latter is non-empty). A general acausal surface on $\tau$ does not necessarily intersect $\gamma^{0}$, but a maximal one will, so the relaxed-minimax surface does as well.
@@
Combining eqns (16) and (17), $|\tilde{\gamma}^{+}|=|\gamma^{+}|$. Further, it follows from the above discussion that $\gamma^{+}$ is the maximal area surface in $\Gamma_{\mathcal{H}}$, and has no more area than any other $\tilde{\gamma}\in\Omega_{\tau}$ on any other time-sheet. Thus, $\gamma^{+}$ is a relaxed-minimax surface. ∎
@@ 3.3 Entanglement wedge nesting & causality of minimax surfaces
We showed in subsec. 2.3 that two minimax surfaces for the same boundary region cannot lie on the same time-sheet. We will now strengthen this statement by showing that they must be fully spacelike-related. We will show this as a consequence of entanglement wedge nesting.
@@ 3.4 Cooperating pairs of time-sheets
The property we show here is that there exists a pair of minimax time-sheets $(\tau_{1},\tau_{2})$ for which the maximal achronal surface on each partial time-sheet is its corresponding partial minimax surface. Note that this property is highly non-trivial: the maximization problem on a partial time-sheet has a boundary condition that depends on the intersection seam $\tau_{1}\cap\tau_{2}$ of both time-sheets, so in general one needs $\tau_{1}$ and $\tau_{2}$ to work together in order for the property to hold. For this reason, we say that when this property holds, the pair of time-sheets $(\tau_{1},\tau_{2})$ is cooperating.888 We postpone a more formal definition to the general case presented in Def. 4.1.
@@
We begin with assuming that $\mathfrak{t}^{\star}\in\mathfrak{M}$, i.e. that $\mathfrak{t}^{\star}$ is minimally intersecting. As a consequence, the two time-sheets intersect forming four partial time-sheets. Let $\mathpzc{i}:=\tau^{\star}_{1}\cap\tau^{\star}_{2}$ be the intersection locus of $\tau^{\star}_{1}$ with $\tau^{\star}_{2}$. In general, $\mathpzc{i}$ may be timelike, spacelike or null (or a piecewise combination of the three). To show (21), assume, for contradiction, that the two pairs of maximal surfaces do not join (i.e. they intersect $\mathpzc{i}$ at different locations). For the moment, we are further going to assume that all of the loci of intersections are different (even between maximal surfaces for different time-sheets). We want to show that if so, then there exists a deformation of $\tau_{1}^{\star}$ and $\tau_{2}^{\star}$ (in the space $\mathfrak{T}$) that would decrease $\max\sum|\gamma_{i}^{(\alpha)}|$, contradicting the minimality in (20). Assume for now that the maximal surfaces $\gamma_{i}^{(\alpha)}$ are unique. [^11] Note that there may be other equal-area minimax surfaces on the full time-sheet. These however will be unstable (or non-achronal in the full spacetime). However, there will be a unique (stable, extremal, achronal) HRT surface by the arguments in 2.3. The set $\mathfrak{T}$ is the set of configurations containing the $\mathscr{J}_{i}$ HRT surfaces, which are then maximal on their respective time-sheets. So when deforming the time-sheet, we do so in a neighborhood of any non-HRT minimax surface, while keeping the HRT surface fixed.
@@
We are now going to prove a theorem that holds for a general pair $\mathscr{J}_{1},\mathscr{J}_{2}$, whose time-sheets are not necessarilty connected, and that ensures that any pair of connected components of a cooperating configuration is minimally intersecting. This defines for us a more general notion of a minimally intersecting configuration, that applies even for disconnected time-sheets. We will see a further generalization of this definition, for arbitrarily many regions, in sec. 4.
@@ 4 The cooperating conjecture
Having introduced the minimax prescription, we can now explore some of its interesting consequences and applications. Contrary to other covariant constructions, here the full spacetime plays a central role. The homology condition, which is usually enforced through a spatial slice, here takes the form of a bulk homology. We begin in subsection 4.1 by presenting the construction of a spacetime graph model for time-dependent states, and explain how a generalization of the cooperating property presented in 3.4 is sufficient for this graph to be equivalent to the RT graph. In 4.2 and 4.3 we explore the validity of this cooperating conjecture — showing that it’s consistent with boundary causality and with the NEC for a class of explicit examples, and comment on some challenging scenarios. We end in 4.4 by providing additional intuition for the cooperating property, explaining how all of the holographic entropy inequalities are geometrically proved with it. The geometric picture makes it clear how minimax resolves the conceptual issues regarding the HRT surfaces not lying on a common bulk slice.
@@ 4.1 A graph model for time-dependent states
As presented in sec. 1, the minimax prescription suggests a way to construct a spacetime graph. Consider a boundary Cauchy slice with $\mathsf{N}$ boundary regions specified. Applying the minimax prescription to all the possible single and joint parties, we obtain a decomposition of the spacetime into connected codimension-0 cells by taking the intersections of all of the homology volumes and their complements. Equivalently, the collection of time-sheets for all subsystems partitions the spacetime into these cells. Each cell corresponds to a vertex in the graph, and vertices corresponding to neighbouring cells are connected with an edge $e$, which is assigned a weight $w(e)$ equal to the area of $\gamma^{+}(e)$, the part of the minimax surface lying on the shared partial time-sheet $\tau(e)$. Cells adjacent to the boundary are called boundary vertices; each one is associated with one elementary region. See fig. 13 for the construction of a spacetime graph for a simple example of non-intersecting time-sheets.
Now that we have defined a spacetime graph, we can compute minimal cuts on it. A cut of a graph is a collection of vertices $U$ that bi-partitions the graph into two sets. To this cut, one associates a cut of edges $\mathscr{C}(U)$ which is the set of all edges that have one end in $U$ and the other end in $U^{c}$; its total weight is the sum of their weights:
@@
A cut $U$ is said to be homologous to a composite subsystem $\mathscr{J}$ (denoted $U\sim\mathscr{J}$) if all the boundary vertices in $U$ are the boundary vertices associated to $\mathscr{J}$. One can interpret $\mathscr{C}(U)$ and $U$ as the graph theoretic analogs of a surface and its corresponding homology region.
The analogy with the static case is striking; however we are still missing a crucial fact to make this connection precise: that the min-cut weight among cuts homologous to $\mathscr{J}$ computes the entropy $S(\mathscr{J})$ of the $\mathscr{J}$ subsystem, i.e. that
$\min_{U\sim\mathscr{J}}\|\mathscr{C}(U)\|\,$ (28)
gives the entropy $S(\mathscr{J})$ of the $\mathscr{J}$ subsystem. As we will show in theorem 4.1, this happens if, for each edge $e$, the edge weight $w(e)$ is given by the maximal area surface on $\tau(e)$, the component of the time-sheet shared by the two neighbouring cells; in other words, if the partial minimax $\gamma^{+}(e)$ is the maximal surface on $\tau(e)$. This is just the cooperating property we discussed in subsection 3.4, only we need it to hold for arbitrarily many time-sheets and boundary regions.
@@
If the time-sheet configuration is cooperating, then for any boundary region $\mathscr{J}$ the graph min cut (28) equals the minimax area $S(\mathscr{J})$.
@@
This is where the cooperating property enters the proof: if each partial minimax surface $\gamma^{+}(e)$ is also the maximal segment on the corresponding partial time-sheet, then the total weight $\|\mathscr{C}(U)\|$ is greater than or equal to that of the maximal achronal surface on $\tau(U)$, since it is being maximized subject to a weaker constraint (each partial time-sheet being achronal, versus their union being achronal):
@@
where the maximum is over achronal surfaces in $\tau(U)$. Since this is true for all cuts $U\sim\mathscr{J}$, the min cut on the graph cannot be smaller than the minimax area:
$\min_{U\sim\mathscr{J}}\|\mathscr{C}(U)\|\geq S(\mathscr{J})\,.$ (33)
Meanwhile, among the set of $\mathscr{J}$-cuts in the correct homology class is the cut $U^{\star}$ corresponding to the minimax time-sheet of $\mathscr{J}$ used for the construction of the graph. Its weight $\|\mathscr{C}(U^{\star})\|$ equals the minimax area, so the bound (33) is saturated. ∎
Figure 13: Spacetime graph inside the regulated spacetime for three boundary regions (with their domains of dependence on the boundary shown in red, green and yellow). Left: We show the full spacetime, the three time-sheets are shown with the respective colors and they end on the joint HRT surface $\hat{\gamma}_{\text{HRT}}(ABC)$ and on $\mathcal{I}$ (shown by the “tent” of null congruences shot from $\hat{\gamma}_{\text{HRT}}(ABC)$). In this simple example the time-sheets do not intersect, and there are only four spacetime cells formed by the union and intersections of the various homology regions: $R(A)$ (yellow node), $R(B)$ (red node), $R(C)$ (green node) and $R^{c}(A)\cap R^{c}(B)\cap R^{c}(C)$ (black node). Right: Same figure but shown from the top for additional clarity.
@@
While minimally intersecting configurations share some properties with collections of RT surfaces, to draw complete analogy to the static case and to allow for the construction of the spacetime graph model we need the (much stronger) cooperating constraint defined 4.1. Our main conjecture is a generalization of theorem 3.5 for an arbitrary configuration:
@@ 4.2.1 Pure AdS3
More precisely, the configuration is constructed as follows – we refer the reader to appendix A for a detailed computation of the spacelike and timelike geodesics. Begin by defining a static rib with opening angle $\Delta\phi$. We now boost the boundary region with some boost parameter $\chi$, so the region’s endpoints have time separation $\Delta t=\Delta\phi\tanh{\chi}$ ). The second rib is constructed in a similar manner, by enlarging the region to have an initial opening angle $\Delta\phi^{\prime}=\Delta\phi+\delta\phi$. We boost the boundary region oppositely by $-\chi$, obtaining $\Delta t^{\prime}=-\Delta\phi^{\prime}\tanh{\chi}$ ). The boost parameter $\chi$ cannot be arbitrarily large as it is constrained by boundary causality; the six endpoints must be mutually spacelike in order to be placed on a common boundary Cauchy slice, see figure 16. We denote the maximum boost allowed by causality with $\chi_{\text{c}}$, and it reads [^18] If the two ribs are near each other, which is the scenario we are considering, this constraint is enforced by demanding the region bounded by the ribs’ endpoints (labeled $A$ in fig. 16) to be spacelike.

### v2

@@ 1 Introduction
For multiple regions, we define the notion of a cooperating time-sheet configuration, as follows: Intersecting time-sheets cut each other up into partial time-sheets, and cut the surfaces they host into partial surfaces; a set of minimax time-sheets cooperates if every partial HRT surface is maximal on its partial time-sheet. The value of cooperating time-sheets is that, as we show, they define a graph model. We also show that, for a pair of crossing regions with connected HRT surfaces, there does exist a cooperating time-sheet pair. However, we have not been able to prove that a cooperating configuration exists for a general set of boundary regions, and our investigations of specific examples are inconclusive; while in some cases we can construct cooperating time-sheets, in others we can actually argue against their existence. This leaves several logically possible outcomes for this story:
@@ 1.1 User-friendly summary
In this section we provide an overview of the main results; proofs and details are given in the main text. In what follows, the bulk spacetime will be denoted by $\hat{\mathcal{M}}$, the conformal boundary by $\hat{\mathcal{N}}$, and the future and past boundaries by $\hat{\mathcal{I}}^{+}$ and $\hat{\mathcal{I}}^{-}$ respectively (we use the hatted notation to avoid confusion with the regulated spacetime that we will use in the bulk of the paper, which will be introduced in subsection 2.1). For a fixed boundary region $A$ on a boundary Cauchy slice $\Sigma\subset\hat{\mathcal{N}}$, the conformal boundary can then be decomposed as (cf. figure 1)
@@
Except for the case of two intersecting time-sheets, we do not have a general proof of the existence of cooperating time-sheets, which we leave as a conjecture. However, we show supporting evidence of its validity in subsection 4.2 by doing direct analytic and numerical computations for a specific set of configurations of time-sheets in pure AdS3 and in asymptotically AdS3 with spherically symmetric NEC-obeying matter. In subsection 4.3, we also present a scenario in pure AdS3 where the existence of a cooperating configuration for three time-sheets is challenged. The setup entails an apparently frustrated configuration of nearly-intersecting but relatively-boosted HRT surfaces. In this extreme scenario cooperation appears difficult to achieve, suggesting that either a more clever construction is needed or the cooperating conjecture, as currently stated, may be false. These potential counterexamples occur in a topologically trivial $2+1$ dimensional bulk, where we know that the RT inequalities are obeyed Czech:2019lps ; thus, while they call into question the cooperating conjecture, they do not directly argue against the equality of the HRT and RT cones. We leave further investigations to future studies.
@@ 2.3 Stability and extremality of minimax surfaces
Of course, these theorems do not preclude a region $A$ having multiple, degenerate minimax/HRT surfaces. However, these surfaces must be maximal on different time-sheets. In subsection 3.3 below, we will prove a stronger result: degenerate minimax surfaces must be fully spacelike-separated. The same result is also obtained from maximin, where it is shown that degenerate HRT surfaces share a common Cauchy slice Wall_2014 ; Grado-White:2024gtx .
@@ 2.4 Relaxed minimax
As noted in subsection 2.2, the definition of time-sheet implies that $\tau$ intersects $\gamma^{0}$ (if the latter is non-empty). A general acausal surface on $\tau$ does not necessarily intersect $\gamma^{0}$, but a maximal one will, so the relaxed-minimax surface does as well.
@@
Combining equations (16) and (17), $|\tilde{\gamma}^{+}|=|\gamma^{+}|$. Further, it follows from the above discussion that $\gamma^{+}$ is the maximal area surface in $\Gamma_{\mathcal{H}}$, and has no more area than any other $\tilde{\gamma}\in\Omega_{\tau}$ on any other time-sheet. Thus, $\gamma^{+}$ is a relaxed-minimax surface. ∎
@@ 3.3 Entanglement wedge nesting & causality of minimax surfaces
We showed in subsection 2.3 that two minimax surfaces for the same boundary region cannot lie on the same time-sheet. We will now strengthen this statement by showing that they must be fully spacelike-related. We will show this as a consequence of entanglement wedge nesting.
@@ 3.4 Cooperating pairs of time-sheets
The property we show here is that there exists a pair of minimax time-sheets $(\tau_{1},\tau_{2})$ for which the maximal achronal surface on each partial time-sheet is its corresponding partial minimax surface. Note that this property is highly non-trivial: the maximization problem on a partial time-sheet has a boundary condition that depends on the intersection seam $\tau_{1}\cap\tau_{2}$ of both time-sheets, so in general one needs $\tau_{1}$ and $\tau_{2}$ to work together in order for the property to hold. For this reason, we say that when this property holds, the pair of time-sheets $(\tau_{1},\tau_{2})$ is cooperating.888 We postpone a more formal definition to the general case presented in definition 4.1.
@@
We begin with assuming that $\mathfrak{t}^{\star}\in\mathfrak{M}$, i.e. that $\mathfrak{t}^{\star}$ is minimally intersecting. As a consequence, the two time-sheets intersect forming four partial time-sheets. Let $\mathpzc{i}:=\tau^{\star}_{1}\cap\tau^{\star}_{2}$ be the intersection locus of $\tau^{\star}_{1}$ with $\tau^{\star}_{2}$. In general, $\mathpzc{i}$ may be timelike, spacelike or null (or a piecewise combination of the three). To show (21), assume, for contradiction, that the two pairs of maximal surfaces do not join (i.e. they intersect $\mathpzc{i}$ at different locations). For the moment, we are further going to assume that all of the loci of intersections are different (even between maximal surfaces for different time-sheets). We want to show that if so, then there exists a deformation of $\tau_{1}^{\star}$ and $\tau_{2}^{\star}$ (in the space $\mathfrak{T}$) that would decrease $\max\sum|\gamma_{i}^{(\alpha)}|$, contradicting the minimality in (20). Assume for now that the maximal surfaces $\gamma_{i}^{(\alpha)}$ are unique. [^11] Note that there may be other equal-area minimax surfaces on the full time-sheet. These however will be unstable (or non-achronal in the full spacetime). However, there will be a unique (stable, extremal, achronal) HRT surface by the arguments in subsection 2.3. The set $\mathfrak{T}$ is the set of configurations containing the $\mathscr{J}_{i}$ HRT surfaces, which are then maximal on their respective time-sheets. So when deforming the time-sheet, we do so in a neighborhood of any non-HRT minimax surface, while keeping the HRT surface fixed.
@@
We are now going to prove a theorem that holds for a general pair $\mathscr{J}_{1},\mathscr{J}_{2}$, whose time-sheets are not necessarilty connected, and that ensures that any pair of connected components of a cooperating configuration is minimally intersecting. This defines for us a more general notion of a minimally intersecting configuration, that applies even for disconnected time-sheets. We will see a further generalization of this definition, for arbitrarily many regions, in section 4.
@@ 4 The cooperating conjecture
Having introduced the minimax prescription, we can now explore some of its interesting consequences and applications. Contrary to other covariant constructions, here the full spacetime plays a central role. The homology condition, which is usually enforced through a spatial slice, here takes the form of a bulk homology. We begin in subsection 4.1 by presenting the construction of a spacetime graph model for time-dependent states, and explain how a generalization of the cooperating property presented in 3.4 is sufficient for this graph to be equivalent to the RT graph. In subsections 4.2 and 4.3 we explore the validity of this cooperating conjecture — showing that it’s consistent with boundary causality and with the NEC for a class of explicit examples, and comment on some challenging scenarios. We end in 4.4 by providing additional intuition for the cooperating property, explaining how all of the holographic entropy inequalities are geometrically proved with it. The geometric picture makes it clear how minimax resolves the conceptual issues regarding the HRT surfaces not lying on a common bulk slice.
@@ 4.1 A graph model for time-dependent states
Figure 13: Spacetime graph inside the regulated spacetime for three boundary regions (with their domains of dependence on the boundary shown in red, green and yellow). Left: We show the full spacetime, the three time-sheets are shown with the respective colors and they end on the joint HRT surface $\hat{\gamma}_{\text{HRT}}(ABC)$ and on $\mathcal{I}$ (shown by the “tent” of null congruences shot from $\hat{\gamma}_{\text{HRT}}(ABC)$). In this simple example the time-sheets do not intersect, and there are only four spacetime cells formed by the union and intersections of the various homology regions: $R(A)$ (yellow node), $R(B)$ (green node), $R(C)$ (red node) and $R^{c}(A)\cap R^{c}(B)\cap R^{c}(C)$ (black node). Right: Same figure but shown from the top for additional clarity.
As presented in section 1, the minimax prescription suggests a way to construct a spacetime graph. Fix a boundary Cauchy slice with $\mathsf{N}$ boundary regions specified. Let $\mathfrak{t}$ be a set of minimax time-sheets for all elementary and composite regions. The time-sheets of $\mathfrak{t}$ partition the bulk into connected codimension-0 cells. We associate to each cell a vertex in the graph. If two cells share a common partial time-sheet, we connect their vertices with an edge $e$, which is assigned a weight $w(e)$ equal to the area of $\gamma^{+}(e)$, the part of the minimax surface lying on the shared partial time-sheet $\tau(e)$. Cells adjacent to the boundary are called boundary vertices; each one is associated with one elementary region. Let $G(\mathfrak{t})$ be the resulting weighted graph. See figure 13 for the construction of a spacetime graph for a simple example with non-intersecting time-sheets.
Now that we have defined a spacetime graph, we can compute minimal cuts on it. A cut $U$ of $G(\mathfrak{t})$ is a subset of its vertices. The cut of edges $\mathscr{C}(U)$ is the set of all edges with one end in $U$ and the other in $U^{c}$; its total weight is the sum of their weights:
@@
A cut $U$ is said to be homologous to a composite region $\mathscr{J}$, denoted $U\sim\mathscr{J}$, if the boundary vertices in $U$ are precisely the boundary vertices associated to $\mathscr{J}$. One can interpret $\mathscr{C}(U)$ and $U$ as the graph-theoretic analogs of a surface and its corresponding homology region respectively. We call the minimal weight for a given region its graph entropy $S_{G(\mathfrak{t})}(\mathscr{J})$:
$S_{G(\mathfrak{t})}(\mathscr{J}):=\min_{U\sim\mathscr{J}}\|\mathscr{C}(U)\|\,.$ (28)
Given that the graph $G(\mathfrak{t})$ is constructed out of minimax time-sheets and surfaces, each of which computes the HRT entropy of a region $\mathscr{J}$, one might expect that the graph entropy equals the HRT entropy. Indeed, the time-sheet for $\tau(\mathscr{J})$ in the configuration $\mathfrak{t}$ corresponds tautologically to a particular cut $U_{\tau(\mathscr{J})}\sim\mathscr{J}$, which simply includes all vertices corresponding to cells in the homology volume for $\tau(\mathscr{J})$. This however only implies a bound,
$S_{G(\mathfrak{t})}(\mathscr{J})\leq S(\mathscr{J})\,,$ (29)
as the graph could include a shortcut: a cut $U\sim\mathscr{J}$ smaller than $U_{\tau(\mathscr{J})}$.
What we will now show is that the cooperating condition on $\mathfrak{t}$ forbids shortcuts, ensuring equality in (29). We already defined the cooperating property for pairs of time-sheets in section 3.4. We will now define it for general sets of time-sheets.
@@
If the time-sheet configuration $\mathfrak{t}$ is cooperating, then for any boundary region $\mathscr{J}$,
$S_{G(\mathfrak{t})}(\mathscr{J})=S(\mathscr{J})\,.$ (30)
@@
This is where the cooperating property enters the proof: if each partial minimax surface $\gamma^{+}(e)$ is also the maximal segment on the corresponding partial time-sheet, then the total weight $\|\mathscr{C}(U)\|$ is greater than or equal to that of the maximal achronal surface on $\tau(U)$, since it is being maximized subject to a weaker constraint (each partial surface being achronal, versus their union being achronal):
@@
where the maximum is over achronal surfaces in $\tau(U)$. Since this is true for all cuts $U\sim\mathscr{J}$, the graph entropy cannot be smaller than the HRT area:
$S_{G(\mathfrak{t})}(\mathscr{J})\geq S(\mathscr{J})\,.$ (35)
Given (29), this implies (30). ∎
@@
While minimally intersecting configurations share some properties with collections of RT surfaces, to draw complete analogy to the static case and to allow for the construction of the spacetime graph model we need the (much stronger) cooperating constraint specified in the definition 4.1. Our main conjecture is a generalization of theorem 3.5 for an arbitrary configuration:
@@ 4.2.1 Pure AdS3
More precisely, the configuration is constructed as follows – we refer the reader to appendix A for a detailed computation of the spacelike and timelike geodesics. Begin by defining a static rib with opening angle $\Delta\phi$. We now boost the boundary region with some boost parameter $\chi$, so the region’s endpoints have time separation $\Delta t=\Delta\phi\tanh{\chi}$ ). The second rib is constructed in a similar manner, by enlarging the region to have an initial opening angle $\Delta\phi^{\prime}=\Delta\phi+\delta\phi$. We boost the boundary region oppositely by $-\chi$, obtaining $\Delta t^{\prime}=-\Delta\phi^{\prime}\tanh{\chi}$ ). The boost parameter $\chi$ cannot be arbitrarily large as it is constrained by boundary causality; the six endpoints must be mutually spacelike in order to be placed on a common boundary Cauchy slice, see figure 16. We denote the maximum boost allowed by causality with $\chi_{\text{c}}$, and it reads [^18] If the two ribs are near each other, which is the scenario we are considering, this constraint is enforced by demanding the region bounded by the ribs’ endpoints (labeled $A$ in figure 16) to be spacelike.